OpenStudy (anonymous):

How would you start this problem? sqrt3sinx+cosx=1

OpenStudy (anonymous):

Ok I did that and I got 3sin^2x+cos^2=1. Then what?

hartnn (hartnn):

or you can divide by 2 on both sides...

hartnn (hartnn):

ohh.. i thought the question was \(\sqrt3 \sin x+\cos x=1\)

OpenStudy (anonymous):

it is

hartnn (hartnn):

then you need to divide by 2 on both sides....

hartnn (hartnn):

and use the identity sin A cos B +cos A sin B =.... ?

OpenStudy (anonymous):

where is the 2 coming from?

hartnn (hartnn):

you have sqrt 3 but for standard angle, sin or cos value is sqrt 3/2 hence , divide by 2

OpenStudy (tkhunny):

I was thinking of using the transformation: \(2\cdot\sin\left(\dfrac{\pi}{6} + x\right) = 1\). I like hartnn's solution better, but unique answers don't care how you find them.

hartnn (hartnn):

i'll get same thing....

hartnn (hartnn):

doubts ? should i give next step ?

OpenStudy (anonymous):

once you divide by 2 what do you get. I am having a hard time figuring out what that looks like?

OpenStudy (mertsj):